OpenStudy (anonymous):

What is the solution of the equation 4x + 17 = 45?

OpenStudy (anonymous):

7 4 x 7 = 28 + 17 = 45

OpenStudy (anonymous):

Just do 45-17 and divide that by 4. Basically you reverse the equation.

OpenStudy (anonymous):

4x+17=45 subtract 17 from both sides 4x=28 Divide by 4 on both sides x=7